OpenStudy (anonymous):

My text seems to prefer -15/a^12 instead of 1/15a^12 I would like to verify that they are the same value, and that I am thinking clearly. Thank you.

OpenStudy (phi):

they are different

OpenStudy (phi):

15^(-1)= 1/15 -15 is -15

OpenStudy (anonymous):

Obviously, thery are not the same expression. The first expression is negative, for all a not 0. The second expression is positive for all a not 0.

OpenStudy (anonymous):

I am asking because I questioned it. It is not obvious to me. Sorry. Should I be asking these questions somewhere else, Easyaspi314?

OpenStudy (anonymous):

Look at the first expression. I see -15 in the numerator and a positive number in the denominator; so a negative number divided by a positive number is negative.

OpenStudy (phi):

\[ \frac{-15}{a^{12}} ≠ \frac{1}{15a^{12}} \]

OpenStudy (anonymous):

Keep in mind that a^12 is always positive, even if a is a negative numbert, because you are raising it to an even power, 12. Of course a cannot be zero. But a can be any number other than zero.

OpenStudy (anonymous):

Thanks all. I understand quite clearly now.

OpenStudy (anonymous):